# From $U(1)\times U(1)$ Symmetry Breaking to Majoron Cosmology: Insights from NANOGrav 15-year Data

### v1

@@ III.2 Gravitational Waves from local cosmic strings
[…]
@@ III.3 Gravitational Waves from cosmic strings in the modified majoron model
The number of maxima in the scalar potential of Eq. (7) gives the number of walls. If $w$ and $w^{\prime}$ are the winding numbers of fields $\phi$ and $\phi^{\prime}$, respectively, then the number of DWs bounded by Type-$\phi^{\prime}$ strings is given by [35, 58] $N_{W}^{\phi^{\prime}}=|nw+n^{\prime}w^{\prime}|_{\text{min}}$, where the minimum is w.r.t varying $w\in\mathbb{Z}$. Also, the energy density stored in a majoron field around a string (known as gradient energy) is proportional to $(nw+n^{\prime}w^{\prime})^{2}$ [35]. Minimal Type-$\phi^{\prime}$ strings, which are expected to be the most abundant, possess $w^{\prime}=\pm 1$ and $w$ is tuned to minimise the gradient energy, which also minimises $N_{W}^{\phi^{\prime}}$. When $\phi$ acquires a VEV, Type-$\phi$ strings form even where Type-$\phi^{\prime}$ strings do not exist, and they have $w^{\prime}=0$. Hence, the number of DWs bounded by Type-$\phi$ strings is given by $N_{W}^{\phi}=|nw|$. For the minimal Type-$\phi$ strings, we have $w=\pm 1$. The ‘domain wall problem’ does not exist if any of the three conditions are satisfied [35, 58]:
@@
In this paper, we explore $n=1$ and $n^{\prime}=1$ conditions only. We leave $N_{W}>1$ DW formation scenarios for a future study, which will require the introduction of further bias term(s) to make the stable DW network dissipate energy and emit GW.
A CS is a topologically stable object, and a conserved topological flux flows along the string. Since Type-$\phi$ and Type-$\phi^{\prime}$ strings have different winding numbers, they form separate string loops, and mixed string loops do not form because mixed string loops do not conserve topological flux. In other words, there cannot be any string loop where one segment is made of a Type-$\phi$ string and another of a Type-$\phi^{\prime}$ string. Hence, Type-$\phi$ and Type-$\phi^{\prime}$ string loops vibrate with separate string tensions and radiate GW.
@@
Similarly the $m_{\chi}=3H(T_{*})$ condition has some impact on Type-$\phi^{\prime}$ LCSs as well, as at temperature $T_{*}$ the whole string-wall network collapses. Hence, the string-wall network should survive to such a temperature $T_{*}$ so that the GW spectrum of LCSs should extend down to $f\sim 10^{-9}$ Hz. The frequency-temperature relationship for LCSs can be found in Eq. (3.5) of Ref. [86], and the relation has a VEV dependence. Using this equation and $m_{\chi}=3H(T_{*})$ we find that for $\eta^{\prime}\sim 10^{14}\,(10^{13})$ GeV, the Type-$\phi^{\prime}$ LCS GW spectrum spans the NANOGrav frequency range for $m_{\chi}\lesssim 10^{-23}\,(10^{-25})$ eV. We get different cut-off values of $m_{\chi}$ for the LCS and GCS-like spectra of Eq. (30) because the $f-T$ relationships are different for them [86, 67].
@@
Another comment is in order for the GW spectrum of the modified majoron model. When the $m_{\chi}=3H(T_{*})$ condition is achieved, DWs will form and assuming that one of the conditions of Eq. (27) is satisfied, the system of DW and strings will collapse. This process will emit additional GWs supplementing that from Eq. (31). However, the spectrum of GWs from the collapse of the Type-$\phi$ CS-DW system is not well known but is expected to have amplitudes and frequencies of the same order as the last e-folding of the scaling regime [87]. Hence, the IR cut-off behaviour of the GW spectrum is expected to be affected. However, if the majoron mass is significantly lower than $10^{-23}$ eV, this contribution is unimportant for our study as the IR cut-off will safely fall below the nano-Hz range. On the other hand, the effect of DW destroying LCSs is known, but for $T_{*}\sim 10^{-7}$ GeV corresponding to $m_{\chi}\sim 10^{-23}$ eV, the deviation from a pure LCS GW spectrum is expected to be minimal [88].
@@
  Total Type-$\phi$ Type-$\phi^{\prime}$
@@ IV Results
Figure 2: Posterior distributions of GCS and SMBHB parameters from comparing the simplified majoron model with the NANOGrav 15-year data in the GCS+SMBHB scenario. The darker (lighter) shade of blue signifies the $1\sigma$ ($2\sigma$) credible contours. Among the four dashed lines of both $\gamma$ as well as $\log_{10}$ the outer (inner) two describe the $2\sigma$ ($1\sigma$) credible intervals. However, for the $\log_{10}(\frac{\eta}{\textrm{GeV}})$ the left (right) one among the two dashed lines signify the upper limits of the $1\sigma$ ($2\sigma$) credible intervals.
@@
In Fig. 1, we present the posterior distributions of the BSM parameter $\log_{10}\big{(}\frac{\eta}{\text{GeV}}\big{)}$, when the GW emission from GCSs is considered the only source of stochastic GW background. The fit to the NANOGrav 15-year data yields a narrow $2\sigma$ Credible Interval of $2.69\times 10^{15}\,\text{GeV}<\eta<3.63\times 10^{15}\,\text{GeV}$. However, when this fit is compared against the reference SMBHB merger GW spectrum, we obtain a tiny BF of $7.66\times 10^{-5}$. That means that the GW emission from the SMBHB merger is obviously a much-preferred solution to the latest NANOGrav data compared to the GCS solution. We can surmise that the result is quite expected because of the negative ($f^{-1/3}$) power law of the GCS GW-spectrum, while NANOGrav data prefers a positive power law. Nevertheless, the above fit can be treated as a constraint on the global $U(1)_{L}$ breaking scale $\eta$ if one assumes GCSs as the only source of stochastic GW. It is also important to recall here that $\eta$ in the simplest majoron model shows up in the expression for light neutrino masses (cf. Eq. (3)). Cosmology places a strong bound on the sum of neutrino masses $\sum m_{\nu}\lesssim 0.1$ eV [90]. Hence, assuming the normal hierarchy (NH) of light neutrino masses and $m_{1}\approx m_{2}\ll m_{3}\sim 0.1$ eV, one can translate the bound on $\eta$ to a simplistic qualitative bound on the effective coupling $16.9712.57<\delta_{33}<$ 16.97 using Eq. (3). As for the inverted hierarchy of neutrino masses, since two of the heavier SM neutrinos are expected to be nearly degenerate, we refrain from deriving such simplistic bounds on any $\delta_{IJ}$ couplings. These results are summarized in Table 2.
@@
  GCS+SMBHB $\log_{10}\big{(}\frac{\eta}{\text{GeV}}\big{)}$ [12,16] $<15.10$ $<5.89$ $0.65\pm 0.16$
@@
  Type-A $\log_{10}\big{(}\frac{\eta^{\prime}}{\text{GeV}}\big{)}$ [11.8,16] [13.67, 14.00] $<0.15$   $0.07\pm 0.01$
$\log_{10}\big{(}\frac{\eta}{\text{GeV}}\big{)}$ [6,16] $<13.50$
  Type-A+SMBHB $\log_{10}\big{(}\frac{\eta^{\prime}}{\text{GeV}}\big{)}$ [11.8,16] [12.46, 13.93] $<0.03$   $0.28\pm 0.05$
$\log_{10}\big{(}\frac{\eta}{\text{GeV}}\big{)}$ [6,16] $<12.85$
@@
  Type-B $\log_{10}\big{(}\frac{\eta^{\prime}}{\text{GeV}}\big{)}$ [11.8,16] [13.58, 13.87] $<0.15$   $0.15\pm 0.02$
$\log_{10}\big{(}\frac{\eta}{\text{GeV}}\big{)}$ [6,16] $<13.50$
  Type-B+SMBHB $\log_{10}\big{(}\frac{\eta^{\prime}}{\text{GeV}}\big{)}$ [11.8,16] $[12.12,13.84]$ $<0.03$   $0.35\pm 0.05$
$\log_{10}\big{(}\frac{\eta}{\text{GeV}}\big{)}$ [6,16] $<12.85$
@@
  Type-C $\log_{10}\big{(}\frac{\eta^{\prime}}{\text{GeV}}\big{)}$ [11.8,16] [13.47, 13.78] $<0.09$   $0.15\pm 0.02$
$\log_{10}\big{(}\frac{\eta}{\text{GeV}}\big{)}$ [6,16] $<13.30$
  Type-C+SMBHB $\log_{10}\big{(}\frac{\eta^{\prime}}{\text{GeV}}\big{)}$ [11.8,16] $[12.10,13.74]$ $<0.03$   $0.28\pm 0.04$
$\log_{10}\big{(}\frac{\eta}{\text{GeV}}\big{)}$ [6,16] $<12.80$
@@
  Type-D $\log_{10}\big{(}\frac{\eta^{\prime}}{\text{GeV}}\big{)}$ [11.8,16] [13.65, 13.95] $<0.15$   $0.12\pm 0.02$
$\log_{10}\big{(}\frac{\eta}{\text{GeV}}\big{)}$ [6,16] $<13.50$
  Type-D+SMBHB $\log_{10}\big{(}\frac{\eta^{\prime}}{\text{GeV}}\big{)}$ [11.8,16] [12.50, 13.88] $<0.04$   $0.42\pm 0.07$
$\log_{10}\big{(}\frac{\eta}{\text{GeV}}\big{)}$ [6,16] $<12.95$
@@
In the modified majoron model, the SM gauge group is extended by a local $U(1)_{B-L}$ and an approximate global $U(1)$ symmetry. When $\phi^{\prime}$ acquires a VEV $(\eta^{\prime})$, it spontaneously breaks $U(1)_{B-L}$. If $\eta^{\prime}>\eta$, the global $U(1)$ remains intact, up to a small soft breaking induced by the operator in Eq. (6). Later, when $\phi$ develops a VEV $(\eta)$, the global $U(1)$ is also spontaneously broken. We assume that one of the conditions of Eq. (27) is satisfied so that only one DW is attached to either Type-$\phi$ or Type-$\phi^{\prime}$ strings. Then, in the scaling regime, the stochastic GW spectrum for this model is given by Eq. (31) for $m_{\chi}<10^{-23}$ eV. The spectrum is a combination of the GW spectrum from LCS (Type-$\phi^{\prime}$ strings) and the GCS-like spectrum (Type-$\phi$ strings), with an IR cutoff regulated by the majoron mass. We consider four different cases, Type-A, Type-B, Type-C, and Type-D, for the GW spectra from this model, depending on the LCS template as described in Table 1 (for more details, please see Sections III.2 and III.3). In this paper, we focus exclusively on the $\eta<\eta^{\prime}$ scenario. The BSM parameters that enter the GW spectrum in Eq. (31) are VEVs $\eta^{\prime}$ and $\eta$. We adopt uniform priors for $\log_{10}\big{(}\frac{\eta^{\prime}}{\text{GeV}}\big{)}$ and $\log_{10}\big{(}\frac{\eta}{\text{GeV}}\big{)}$ in the range [11.8, 16] and [6, 16], respectively 222 We do not attempt to lower the prior range for the VEV $\eta^{\prime}$ for Type-$\phi^{\prime}$ LCS. This limitation stems from using the PTArcade LCS templates, which are valid in the range of $[-14,0]$ for $\log_{10}(\text{G}\mu)$—equivalent to $\log_{10}\big{(}\frac{\eta^{\prime}}{\text{GeV}}\big{)}$ spanning from 11.83 to 18.83. .
@@
Since the GW spectrum of Eq. (30) is not significantly different from that of pure GCS, we expect results similar to those of the simplified majoron model. In the Type-$\phi$ strings only scenario, the fit to the NANOGrav data returns a $2\sigma$ Credible Interval of $2.75\times 10^{15}\,\text{GeV}<\eta<3.63\times 10^{15}\,\text{GeV}$, which is almost identical to the GCS-only scenario in the simplified majoron model. When Type-$\phi$ strings are considered as a secondary source of GW emission on top of SMBHB mergers, we obtain $2\sigma$ Credible Intervals of $\eta<1.26\times 10^{15}$ GeV, $-14.77<\log_{10}A<-14.24$, and $4.703.60<\gamma<$ 4.70, respectively. Again, the results are similar to those of the GCS+SMBHB scenario in the simplified majoron model. Thus, we do not show any separate plots for $10^{-23}\,\text{eV}<m_{\chi}<10^{-15}\,\text{eV}$ regime in the modified majoron model.

### v2

@@ III.2 Gravitational Waves from local cosmic strings
$\Omega_{\text{GW}}^{(k)}(f)\Bigg|_{\text{LCS}}=\frac{\mathcal{D}_{\alpha}}{\rho_{c}}\frac{2k}{f}\frac{F_{\alpha}}{\alpha}\int_{t_{F}}^{t_{0}}d\tilde{t}\frac{\Gamma^{(k)}G\mu^{2}}{\left(\alpha+\Gamma G\mu\right)}\frac{\varepsilon_{\text{loop}}\left(t_{i}^{(k)}\right)}{{t_{i}^{(k)}}^{4}}\theta(\tilde{l})\theta(\tilde{t}-t_{i}){\left(\frac{a(\tilde{t})}{a(t_{0})}\right)}^{5}{\left(\frac{a(t_{i}^{(k)})}{a(\tilde{t})}\right)}^{3}\,$, (25)
@@ III.3 Gravitational Waves from cosmic strings in the modified majoron model
The number of maxima in the scalar potential of Eq. (7) gives the number of walls. If $w$ and $w^{\prime}$ are the winding numbers of fields $\phi$ and $\phi^{\prime}$, respectively, then the number of DWs bounded by Type-$\phi^{\prime}$ strings is given by [39, 62] $N_{W}^{\phi^{\prime}}=|nw+n^{\prime}w^{\prime}|_{\text{min}}$, where the minimum is w.r.t varying $w\in\mathbb{Z}$. Also, the energy density stored in a majoron field around a string (known as gradient energy) is proportional to $(nw+n^{\prime}w^{\prime})^{2}$ [39, 91]. Minimal Type-$\phi^{\prime}$ strings, which are expected to be the most abundant, possess $w^{\prime}=\pm 1$ and $w$ is tuned to minimise the gradient energy, which also minimises $N_{W}^{\phi^{\prime}}$. When $\phi$ acquires a VEV, Type-$\phi$ strings form even where Type-$\phi^{\prime}$ strings do not exist, and they have $w^{\prime}=0$. Hence, the number of DWs bounded by Type-$\phi$ strings is given by $N_{W}^{\phi}=|nw|$. For the minimal Type-$\phi$ strings, we have $w=\pm 1$. The ‘domain wall problem’ does not exist if any of the three conditions are satisfied [39, 62]:
@@
In this paper, we explore only the scenario corresponding to $n=1\,(n^{\prime}=d-1)$. We leave $N_{W}>1$ DW formation scenarios for a future study, which will require the introduction of further bias term(s) to make the stable DW network dissipate energy and emit GW. The details of the benchmark parameters, used to describe the scenarios, are summarized in Table 1 and their phenomenological aspects are outlined later in Table 5.
Benchmarks $d$ $n$ $n^{\prime}$ Type-$\phi$ strings Type-$\phi^{\prime}$ strings
$w$ $w^{\prime}$ $N_{W}^{\phi}=|nw|$ $w$ $w^{\prime}$ $N_{W}^{\phi^{\prime}}=|nw+n^{\prime}w^{\prime}|_{\text{min}}$
BP1 16 1 15 1 0 1 -15 1 0
BP2 17 1 16 1 0 1 -16 1 0
Table 1: Benchmark points (motivated by the relic abundance of the dark matter) showing the values of $d,n,n^{\prime}$ and the wingding numbers $(w,w^{\prime})$ for both Type-$\phi$ and Type-$\phi^{\prime}$ strings, together with the corresponding numbers of domain walls attached to each type. The table highlights only the cases in which $N_{W}^{\phi}=1$ and $N_{W}^{\phi^{\prime}}=0$.
A CS is a topologically stable object, and a conserved topological flux flows along the string. Since Type-$\phi$ and Type-$\phi^{\prime}$ strings have different winding numbers, they form separate string loops, and mixed string loops do not form because mixed string loops do not conserve topological flux. In other words, there cannot be any string loop where one segment is made of a Type-$\phi$ string and another of a Type-$\phi^{\prime}$ string. Also, we assume Type-$\phi$ and Type-$\phi^{\prime}$ strings’ evolution to proceed independently into distinct scaling regimes, with no interaction between them. Hence, Type-$\phi$ and Type-$\phi^{\prime}$ string loops vibrate with separate string tensions and radiate GW.
@@
Similarly the $m_{\chi}=3H(T_{*})$ condition has some impact on Type-$\phi^{\prime}$ LCSs as well, as at temperature $T_{*}$ the whole string-wall network collapses. Hence, the string-wall network should survive to such a temperature $T_{*}$ so that the GW spectrum of LCSs should extend down to $f\sim 10^{-9}$ Hz. The frequency-temperature relationship for LCSs can be found in Eq. (3.5) of Ref. [92], and the relation has a VEV dependence. Using this equation and $m_{\chi}=3H(T_{*})$ we find that for $\eta^{\prime}\sim 10^{14}\,(10^{13})$ GeV, the Type-$\phi^{\prime}$ LCS GW spectrum spans the NANOGrav frequency range for $m_{\chi}\lesssim 10^{-23}\,(10^{-25})$ eV. We get different cut-off values of $m_{\chi}$ for the LCS and GCS-like spectra of Eq. (30) because the fTf-T relationships are different for them [92, 72].
@@
Another comment is in order for the GW spectrum of the modified majoron model. When the $m_{\chi}=3H(T_{*})$ condition is achieved, DWs will form and assuming that one of the conditions of Eq. (27) is satisfied, the system of DW and strings will collapse. This process will emit additional GWs supplementing that from Eq. (31). However, the spectrum of GWs from the collapse of the Type-$\phi$ CS-DW system is not well known but is expected to have amplitudes and frequencies of the same order as the last e-folding of the scaling regime [93]. Hence, the IR cut-off behaviour of the GW spectrum is expected to be affected. However, if the majoron mass is significantly lower than $10^{-23}$ eV, this contribution is unimportant for our study as the IR cut-off will safely fall below the nano-Hz range. On the other hand, the effect of DWs destroying LCSs is known, but for $T_{*}\sim 10^{-7}$ GeV corresponding to $m_{\chi}\sim 10^{-23}$ eV, the deviation from a pure LCS GW spectrum is expected to be minimal [94].
@@
  Total $\phi Type$-$\phi^{\prime}$
@@ IV Results
Figure 2: Posterior distributions of GCS and SMBHB parameters from comparing the simplified majoron model with the NANOGrav 15-year data in the GCS+SMBHB scenario. The darker (lighter) shade of blue signifies the $1\sigma$ ($2\sigma$) credible contours. Among the four dashed lines of both $\gamma$ as well as $\log_{10}$A the outer (inner) two describe the $2\sigma$ ($1\sigma$) credible intervals. However, for the $\log_{10}(\frac{\eta}{\textrm{GeV}})$ the left (right) one among the two dashed lines signify the upper limits of the $1\sigma$ ($2\sigma$) credible intervals.
@@
In Fig. 1, we present the posterior distributions of the BSM parameter $\log_{10}\big(\frac{\eta}{\text{GeV}}\big)$, when the GW emission from GCSs is considered the only source of stochastic GW background. The fit to the NANOGrav 15-year data yields a narrow $2\sigma$ Credible Interval of $2.69\times 10^{15}\,\text{GeV}<\eta<3.63\times 10^{15}\,\text{GeV}$. However, when this fit is compared against the reference SMBHB merger GW spectrum, we obtain a tiny BF of $7.66\times 10^{-5}$. That means that the GW emission from the SMBHB merger is obviously a much-preferred solution to the latest NANOGrav data compared to the GCS solution. We can surmise that the result is quite expected because of the negative ($f^{-1/3}$) power law of the GCS GW-spectrum, while NANOGrav data prefers a positive power law. Nevertheless, the above fit can be treated as a constraint on the global $U(1)_{L}$ breaking scale $\eta$ if one assumes GCSs as the only source of stochastic GW. It is also important to recall here that $\eta$ in the simplest majoron model shows up in the expression for light neutrino masses (cf. Eq. (3)). Cosmology places a strong bound on the sum of neutrino masses $\sum m_{\nu}\lesssim 0.1$ eV [96]. Hence, assuming the normal hierarchy (NH) of light neutrino masses and $m_{1}\approx m_{2}\ll m_{3}\sim 0.1$ eV, one can translate the bound on $\eta$ to a simplistic qualitative bound on the effective coupling $12.57<\delta_{33}<16.97$ using Eq. (3). As for the inverted hierarchy of neutrino masses, since two of the heavier SM neutrinos are expected to be nearly degenerate, we refrain from deriving such simplistic bounds on any $\delta_{IJ}$ couplings. These results are summarized in Table 3.
@@
  GCS+SMBHB $\log_{10}\big(\frac{\eta}{\text{GeV}}\big)$ [12,16] <15.10<15.10 <5.89<5.89 $0.65\pm 0.16$
@@
  Type-A $\log_{10}\big(\frac{\eta^{\prime}}{\text{GeV}}\big)$ [11.8,16] [13.67, 14.00] <0.15<0.15   $0.07\pm 0.01$
$\log_{10}\big(\frac{\eta}{\text{GeV}}\big)$ [6,16] <13.50<13.50
  Type-A+SMBHB $\log_{10}\big(\frac{\eta^{\prime}}{\text{GeV}}\big)$ [11.8,16] [12.46, 13.93] <0.03<0.03   $0.28\pm 0.05$
$\log_{10}\big(\frac{\eta}{\text{GeV}}\big)$ [6,16] <12.85<12.85
@@
  Type-B $\log_{10}\big(\frac{\eta^{\prime}}{\text{GeV}}\big)$ [11.8,16] [13.58, 13.87] <0.15<0.15   $0.15\pm 0.02$
$\log_{10}\big(\frac{\eta}{\text{GeV}}\big)$ [6,16] <13.50<13.50
  Type-B+SMBHB $\log_{10}\big(\frac{\eta^{\prime}}{\text{GeV}}\big)$ [11.8,16] $[12.12,13.84]$ <0.03<0.03   $0.35\pm 0.05$
$\log_{10}\big(\frac{\eta}{\text{GeV}}\big)$ [6,16] <12.85<12.85
@@
  Type-C $\log_{10}\big(\frac{\eta^{\prime}}{\text{GeV}}\big)$ [11.8,16] [13.47, 13.78] <0.09<0.09   $0.15\pm 0.02$
$\log_{10}\big(\frac{\eta}{\text{GeV}}\big)$ [6,16] <13.30<13.30
  Type-C+SMBHB $\log_{10}\big(\frac{\eta^{\prime}}{\text{GeV}}\big)$ [11.8,16] $[12.10,13.74]$ <0.03<0.03   $0.28\pm 0.04$
$\log_{10}\big(\frac{\eta}{\text{GeV}}\big)$ [6,16] <12.80<12.80
@@
  Type-D $\log_{10}\big(\frac{\eta^{\prime}}{\text{GeV}}\big)$ [11.8,16] [13.65, 13.95] <0.15<0.15   $0.12\pm 0.02$
$\log_{10}\big(\frac{\eta}{\text{GeV}}\big)$ [6,16] <13.50<13.50
  Type-D+SMBHB $\log_{10}\big(\frac{\eta^{\prime}}{\text{GeV}}\big)$ [11.8,16] [12.50, 13.88] <0.04<0.04   $0.42\pm 0.07$
$\log_{10}\big(\frac{\eta}{\text{GeV}}\big)$ [6,16] <12.95<12.95
@@
In the modified majoron model, the SM gauge group is extended by a local $U(1)_{B-L}$ and an approximate global $U(1)$ symmetry. When $\phi^{\prime}$ acquires a VEV $(\eta^{\prime})$, it spontaneously breaks $U(1)_{B-L}$. If $\eta^{\prime}>\eta$, the global $U(1)$ remains intact, up to a small soft breaking induced by the operator in Eq. (6). Later, when $\phi$ develops a VEV $(\eta)$, the global $U(1)$ is also spontaneously broken. We assume that one of the conditions of Eq. (27) is satisfied so that only one DW is attached to either Type-$\phi$ or Type-$\phi^{\prime}$ strings. Then, in the scaling regime, the stochastic GW spectrum for this model is given by Eq. (31) for $m_{\chi}<10^{-23}$ eV. The spectrum is a combination of the GW spectrum from LCS (Type-$\phi^{\prime}$ strings) and the GCS-like spectrum (Type-$\phi$ strings), with an IR cutoff regulated by the majoron mass. We consider four different cases, Type-A, Type-B, Type-C, and Type-D, for the GW spectra from this model, depending on the LCS template as described in Table 2 (for more details, please see Sections III.2 and III.3). In this paper, we focus exclusively on the $\eta<\eta^{\prime}$ scenario. The BSM parameters that enter the GW spectrum in Eq. (31) are VEVs $\eta^{\prime}$ and $\eta$. We adopt uniform priors for $\log_{10}\big(\frac{\eta^{\prime}}{\text{GeV}}\big)$ and $\log_{10}\big(\frac{\eta}{\text{GeV}}\big)$ in the range [11.8, 16] and [6, 16], respectively  [^3] We do not attempt to lower the prior range for the VEV $\eta^{\prime}$ for Type-$\phi^{\prime}$ LCS. This limitation stems from using the PTArcade LCS templates, which are valid in the range of [14,0][-14,0] for $\log_{10}(\text{G}\mu)$—equivalent to $\log_{10}\big(\frac{\eta^{\prime}}{\text{GeV}}\big)$ spanning from 11.83 to 18.83. .
@@
Since the GW spectrum of Eq. (30) is not significantly different from that of pure GCS, we expect results similar to those of the simplified majoron model. In the Type-$\phi$ strings only scenario, the fit to the NANOGrav data returns a $2\sigma$ Credible Interval of $2.75\times 10^{15}\,\text{GeV}<\eta<3.63\times 10^{15}\,\text{GeV}$, which is almost identical to the GCS-only scenario in the simplified majoron model. When Type-$\phi$ strings are considered as a secondary source of GW emission on top of SMBHB mergers, we obtain $2\sigma$ Credible Intervals of $\eta<1.26\times 10^{15}$ GeV, $14.77<\log_{10}A<-14.24$, and $3.60<\gamma<4.70$, respectively. Again, the results are similar to those of the GCS+SMBHB scenario in the simplified majoron model. Thus, we do not show any separate plots for $10^{-23}\,\text{eV}<m_{\chi}<10^{-15}\,\text{eV}$ regime in the modified majoron model.